OpenStudy (anonymous):

98 - 4 squared divided by 2 I tried to solve and got the answer 41 but my teacher says that is wrong

OpenStudy (anonymous):

Doe sit look like this: \[(98-4^2)/2\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no there are no parenthesis

OpenStudy (anonymous):

is the answer 90

OpenStudy (anonymous):

First what does 4^2=?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

hmmm I got 41 tooo are you sure that's how it is written?

OpenStudy (anonymous):

sorry it is without the parenthesis

OpenStudy (anonymous):

well all the parenthisis are saying is that both those numbers are on top

OpenStudy (anonymous):

Is it: \[\frac{98−4^2}{2} \text{ or }98-\frac{4^2}{2}\]

OpenStudy (anonymous):

i got 41 , -4^2 =-16, (-4)^2=16 , so i think u state the question wrong